# Three-dimensional Imaging of Pion using Lattice QCD: Generalized Parton Distributions

### v1

@@ III Numerical results on the valence pion light-cone GPDs
Figure 7: The dependence of the valence light-cone GPDs on $z_{s}$ is illustrated at $P_{z}=1.937$ GeV and $-t=0$ GeV2. The black band displays the NLO global analysis results [84]. The red band represents the valence pion PDFs obtained with NNLO matching in a hybrid scheme, excluding LRR and RGR [70].
III.2 The $t$-dependence of the light-cone GPDs
Having analyzed the systematic effects in the determination of the valence pion light-cone GPDs at some representative values of $-t$, we can now study the entire $t$ dependence of the light-cone GPDs in more detail. In Fig. 8, we present the valence light-cone GPDs as a function of the Bjorken-$x$ across all the available values of $-t$. As expected from the previous discussion, for $x<0.2$, there are significant differences in all results between the cases with (solid-filled bands) and without (line-filled bands) RGR. The distributions exhibit significant dependence on $-t$ and $x$. Firstly, the light-cone GPDs $H(x,\xi=0,t)$ decrease as $-t$ increases for a fixed $x$. Secondly, the fall-off of the GPDs along $x$ is notably reduced at larger values of $-t$. Similar behavior of the pion GPDs was also found in Ref. [32], where a lattice spacing of $a=0.09$ fm and a symmetric frame were utilized.
@@ III.2 The $t$-dependence of the light-cone GPDs
To further explore the $t$ and $x$ dependence of the pion valence light-cone GPDs, we would like to parametrize our results with a suitable ansatz. Based on past experience with the pion form factor [71], we employ the monopole form
@@
where $H(x,0)$ represents the GPDs at zero momentum transfer, equivalent to the PDFs, and $M(x)$ is a $x$-dependent monopole mass parameter, which is related to the pion effective radius as $\langle r^{2}(x)\rangle=6/M^{2}(x)$, in analogy to the usual pion charge radius. This fit ansatz works very well, as demonstrated by the bands in the left panel of Fig. 9. The left panel displays the $t$ dependence of the light-cone GPDs for four selected values of $x$, namely $x=0.2,~{}0.3,~{}0.5$, and 0.8. The lattice results are shown as the data points, and the corresponding monopole fit results are displayed as the bands. We can observe that the $t$ dependence of the valence light-cone GPDs becomes milder as $x$ increases. For $x=0.8$, almost no $t$ dependence can be discerned within the estimated errors. The results of the pion effective radius are shown in the right panel of Fig. 9. The effective radius clearly decreases with increasing $x$. This means that when quarks have higher momentum fractions, they are likely to be confined to a smaller spatial region in the transverse plane. This confinement results in narrower distributions in position space, potentially reducing the effective radius. Interestingly, the effective radius for $x=0.2$ is comparable to the pion charge radius obtained from the pion electromagnetic form factor, specifically $\langle r_{\pi}^{2}\rangle=0.313(27)$ fm2, as calculated in Ref. [71] using the same pion mass.
@@
where $\mathbf{b}_{\perp}$ is the impact parameter. The IPDs describe the probability density of finding a parton with momentum fraction $x$ at a specific transverse distance $\mathbf{b}_{\perp}$ from the center of the transverse momentum (CoTM). They can offer a comprehensive view of the parton distributions in both momentum and coordinate spaces within the hadron. Using our monopole fit results of $H(x,t)$, we can perform a Fourier transform in $\mathbf{\Delta}_{\perp}$ to obtain the valence IPDs. Fig. 10 shows the results with varying $x$ from 0.2 to 0.9 and $\mathbf{b}_{\perp}$ from 0 to 1 fm. We can find that when the partons carry a larger momentum fraction $x$, the distributions are more localized, which results in a more concentrated distribution in the impact-parameter space. Conversely, as $x$ decreases, the distributions become broader in $\mathbf{b}_{\perp}$, indicating a more diffuse spatial distribution of lower-momentum quarks.
It is interesting to compare the pion valence IPDs with those of the proton, as shown in Fig. 11. We display the comparison in a two-dimensional distribution at $x=0.3$ and 0.5. The light-cone proton GPDs data are from Ref. [85]. We perform a dipole fit on the proton GPDs and then get the corresponding IPDs through the Fourier transform. For both the pion and proton, the distributions become more concentrated at larger $x$, corresponding to smaller effective radii, which is consistent with our findings from the right panel of Fig. 9. Comparison of the proton and pion results at fixed $x$ reveals that the proton exhibits a broader distribution of valence quarks than the pion at both $x=0.3$ and 0.5. To some extent, this observation relates to the fact that the charge radius of the proton is larger than that of the pion. Fig. 11 also indicates that the valence quark distributions within the pion are somewhat less peaked at the very center compared to those within the proton.
@@
Figure 11: The two-dimensional distributions show a comparison of the IPDs between the pion and proton [85] at $x$=0.3 and 0.5.
IV Conclusions
@@
We study the pion valence light-cone GPDs using the LaMET approach at a lattice spacing of 0.04 fm, which is more than two times smaller than the lattice spacing used in other studies of pion GPDs ($a=0.09$ fm) [32]. In this work, we utilize the Lorentz-invariant definition of the GPDs, enabling us to perform calculations in the non-Breit frame. This approach enables simultaneous calculation of multiple momentum transfers and thus significantly reduces the computational costs. In this work, we perform calculations within both the Breit and non-Breit frames. By comparing the amplitude results of these two frames, we demonstrated once again the validity and efficacy of the Lorentz-invariant approach. To obtain the pion valence light-cone GPDs from the qGPDs, we use NNLO hybrid-scheme matching along with leading renormalon resummation and renormalization group resummation. The use of leading renormalon resummation significantly reduces the perturbative uncertainty in the hybrid renormalization scheme. Furthermore, the use of renormalization group resummation allows us to determine the range of Bjorken-$x$ for which the LaMET approach is reliable for a specific momentum value $P_{z}$. Our results at zero momentum transfer, specifically the valence light-cone PDF results, are in good agreement with the global analyses [84] within the range $0.2\lesssim x\lesssim 0.8$. The $P_{z}$ dependence of the final light-cone GPD results is significantly reduced compared to the qGPDs, which indicates the effectiveness of the perturbative matching framework. Our lattice calculations provide a comprehensive three-dimensional imaging of the pion structure, which can be better visualized in terms of impact-parameter-dependent distribution. From these findings, we observe that the effective transverse size of the pion reduces as $x$ increases, a pattern also observed in the proton [85]. Additionally, the effective size of the pion is consistently smaller than that of the proton for the considered $x$ region.
@@ Acknowledgements.
We thank Shohini Bhattacharya, Yushan Su, and Rui Zhang for their valuable communications. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract Nos. DE-SC0012704, DE-AC02-06CH11357, and within the frameworks of Scientific Discovery through Advanced Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond, and under the umbrella of the Quark-Gluon Tomography (QGT) Topical Collaboration with Award DE-SC0023646. HTD is supported by the NSFC under grant No. 12325508 and the National Key Research and Development Program of China under Grant No. 2022YFA1604900. SS is supported by the National Science Foundation under CAREER Award PHY-1847893. YZ is partially supported by the 2023 Physical Sciences and Engineering (PSE) Early Investigator Named Award program at Argonne National Laboratory. This research used awards of computer time provided by the U.S. Department of Energy’s INCITE and ALCC programs at the Argonne and the Oak Ridge Leadership Computing Facilities. The Argonne Leadership Computing Facility at Argonne National Laboratory is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC02-06CH11357. The Oak Ridge Leadership Computing Facility at Oak Ridge National Laboratory is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC05-00OR22725. This research also used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. Computations for this work were carried out in part on facilities of the USQCD Collaboration, funded by the Office of Science of the U.S. Department of Energy.
@@ Appendix A Fit results of the two-point and three-point correlation functions
In this appendix, we present some detailed analysis of the pion two-point and three-point correlation functions.
As mentioned in the main text, by employing the spectral decomposition formula along with 2-state or 3-state fits of the pion two-point correlators [62, 70, 71, 72, 73], we extract energy values of both the ground and excited states of the pion, as well as the corresponding amplitudes. We show the results of the ground and the first excited state energy for different momenta in Fig. 12. We can see that the fit results for $E_{0}$ agree well with the dispersion relation shown by the solid lines, and in most cases, the results for $E_{1}$ also agree with them within errors.
@@
To obtain the matrix elements, we consider the ratio $R^{\mu}$ given by Eq. (12) and perform 2-state or 3-state fits. Here, we use the energy levels and amplitudes obtained from the corresponding fits of the two-point functions. In Fig. 13, we select two specific cases from the data of $z=0$ and $\mu=t$ to show the results of the ratio as $R^{t}/\bar{P}^{t}$. The left one belongs to the case of $P_{z}$ = 0.97 GeV and $-t$ = 0.95 GeV2, while the right one corresponds to the case of $P_{z}$ = 1.94 GeV and $-t$ = 1.69 GeV2, that is, the largest momentum transfer in this work. The lattice results are shown as the data points with different colors representing the different time separations. The bands with corresponding colors denote the two-state fit results of the lattice data. And the grey bands are the extrapolated results of the ratio, that is, the bare matrix elements. We collect all the bare matrix element results of the largest momentum in Fig. 14. It is evident that the results display a decreasing trend with increasing $z$ and also as the momentum transfer increases, as expected.
Figure 13: Two specific examples of the ratio $R^{t}(z=0)/\bar{P}^{t}$ are shown. Left: $P_{z}$ = 0.97 GeV and $-t$ = 0.95 GeV2, right: $P_{z}$ = 1.94 GeV and $-t$ = 1.69 GeV2. The data points represent the lattice results, and the bands with corresponding colors are the fit results of the lattice data. The grey bands show the results of the bare matrix elements.
@@
Appendix B Perturbative coefficient
B.1 Matching coefficient

### v2

@@ 3 Numerical results on the valence pion light-cone GPDs
Figure 7: The dependence of the valence light-cone GPDs on $z_{s}$ is illustrated at $P_{z}=1.937$ GeV and $-t=0$ GeV2. The black band displays the NLO global analyses results Barry_2021 . The red band represents the valence pion PDFs obtained with NNLO matching in a hybrid scheme, excluding LRR and RGR Gao:2021dbh .
If one uses the hybrid scheme to perform the renormalization, the resulting light-cone GPDs may depend on the choice of $z_{s}$. Therefore, it is important to check the dependence of our results on $z_{s}$. We compare the valence light-cone GPD results with three different values of $z_{s}$ for $P_{z}=1.937$ GeV in figure 7. As a demonstration, we show our results for $-t=0$ GeV2, obtained using NNLO+LRR+RGR. The lines represent the central values obtained with $\kappa=1.414$, while the bands encompass both statistical and systematic errors. This convention for the error bands is used in the subsequent figures as well. We can see that the final valence light-cone GPD results have little dependence on $z_{s}$ for $x>0.2$. At $0.30.2<x<$ 0.3, the results at $z_{s}=2a$ show a slight deviation from the other two cases, which is probably caused by the discretization effects as $z_{s}\sim a$. The situations are similar for other values of $-t$. Thus, our predictions in the region $0.2\lesssim x\lesssim 0.8$ are not sensitive to the choice of $z_{s}$. Considering that the value of $\bar{m}_{0}$ at $z/a=3$ exhibits the minimal dependence on the perturbative order and renormalization scale, we adopt the corresponding position $z_{s}/a=3$ for subsequent analyses.
Furthermore, the zero momentum transfer case corresponds to the PDF results, allowing for comparison with previous PDF calculations. Specifically, we compare our findings with the global analyses at NLO accuracy (JAM21) Barry_2021 and our previous results using the NNLO matching (without LRR or RGR) within the same lattice framework (BNL-ANL21) Gao:2021dbh . Our results agree well with the global fit results for $0.2\lesssim x\lesssim 0.8$. Regarding the comparison with BNL-ANL21, the results show good agreement in the range $0.3\lesssim x\lesssim 0.7$, while the differences observed in the small and large $x$ regions are expected due to contributions from both LRR and RGR.
3.2 The $t$-dependence of the light-cone GPDs
@@
Having analyzed the systematic effects in the determination of the valence pion light-cone GPDs at some representative values of $-t$, we can now study the entire $t$ dependence of the light-cone GPDs in more detail. In figure 8, we present the valence light-cone GPDs as a function of the Bjorken-$x$ across all the available values of $-t$. As expected from the previous discussion, for $x<0.2$, there are significant differences in all results between the cases with (solid-filled bands) and without (line-filled bands) RGR. The distributions exhibit significant dependence on $-t$ and $x$. Firstly, the light-cone GPDs $H(x,\xi=0,t)$ decrease as $-t$ increases for a fixed $x$. Secondly, the fall-off of the GPDs along $x$ is notably reduced at larger values of $-t$. Similar behavior of the pion GPDs was also found in ref. Lin:2023gxz , where a lattice spacing of $a=0.09$ fm and a symmetric frame were utilized.
@@
To further explore the $t$ and $x$ dependence of the pion valence light-cone GPDs, we would like to parametrize our results with a suitable ansatz. Based on past experience with the pion form factor Gao:2021xsm , we employ the monopole form
@@
where $H(x,0)$ represents the GPDs at zero momentum transfer, equivalent to the PDFs, and $M(x)$ is a $x$-dependent monopole mass parameter, which is related to the pion effective radius as $\langle r^{2}(x)\rangle=6/M^{2}(x)$, in analogy to the usual pion charge radius. This fit ansatz works very well, as demonstrated by the bands in the left panel of figure 9. The left panel displays the $t$ dependence of the light-cone GPDs for four selected values of $x\in[0.2,0.3,0.5,0.8]$. The lattice results are shown as the data points, and the corresponding monopole fit results are displayed as the bands. We can observe that the $t$ dependence of the valence light-cone GPDs becomes milder as $x$ increases. For $x=0.8$, almost no $t$ dependence can be discerned within the estimated errors. The results of the pion effective radius are shown in the right panel of figure 9. The asymmetric error bars reflect the underlying skewed distribution of the effective radius sample results. Given that radius cannot be negative, we use the median as the central value, with the lower and upper uncertainties defined by the 16th and 84th percentiles of the sample distribution, respectively. The effective radius clearly decreases with increasing $x$. This means that when quarks have higher momentum fractions, they are likely to be confined to a smaller spatial region in the transverse plane. This confinement results in narrower distributions in position space, potentially reducing the effective radius. Interestingly, the effective radius for $x=0.2$ is comparable to the pion charge radius obtained from the pion electromagnetic form factor, specifically $\langle r_{\pi}^{2}\rangle=0.313(27)$ fm2, as calculated in ref. Gao:2021xsm using the same pion mass.
@@
where $\mathbf{b}_{\perp}$ is the impact parameter. The IPDs describe the probability density of finding a parton with momentum fraction $x$ at a specific transverse distance $\mathbf{b}_{\perp}$ from the center of the transverse momentum (CoTM). They can offer a comprehensive view of the parton distributions in both momentum and coordinate spaces within the hadron. Using our monopole fit results of $H(x,t)$, we can perform a Fourier transform in $\mathbf{\Delta}_{\perp}$ to obtain the valence IPDs. Figure 10 shows the results with varying $x$ from 0.2 to 0.8 and $\mathbf{b}_{\perp}$ from 0 to 1 fm. We can find that when the partons carry a larger momentum fraction $x$, the distributions are more localized, which results in a more concentrated distribution in the impact-parameter space. Conversely, as $x$ decreases, the distributions become broader in $\mathbf{b}_{\perp}$, indicating a more diffuse spatial distribution of lower-momentum quarks.
@@
Figure 11: The two-dimensional distributions show a comparison of the IPDs between the pion and proton Cichy:2023dgk at $x$=0.3 and 0.5.
It is interesting to compare the pion valence IPDs with those of the proton, as shown in figure 11. We display the comparison in a two-dimensional distribution at $x=0.3$ and 0.5. The light-cone proton GPDs data are from ref. Cichy:2023dgk . We perform a dipole fit on the proton GPDs and then get the corresponding IPDs through the Fourier transform. For both the pion and proton, the distributions become more concentrated at larger $x$, corresponding to smaller effective radii, which is consistent with our findings from the right panel of figure 9. Comparison of the proton and pion results at fixed $x$ reveals that the proton exhibits a broader distribution of valence quarks than the pion at both $x=0.3$ and 0.5. To some extent, this observation relates to the fact that the charge radius of the proton is larger than that of the pion. Figure 11 also indicates that the valence quark distributions within the pion are somewhat less peaked at the very center compared to those within the proton.
4 Conclusions
@@
We study the pion valence light-cone GPDs using the LaMET approach at a lattice spacing of 0.04 fm, which is more than two times smaller than the lattice spacing used in other studies of pion GPDs ($a=0.09$ fm) Lin:2023gxz . In this work, we utilize the Lorentz-invariant definition of the GPDs, enabling us to perform calculations in the non-Breit frame. This approach enables simultaneous calculation of multiple momentum transfers and thus significantly reduces the computational costs. In this work, we perform calculations within both the Breit and non-Breit frames. By comparing the amplitude results of these two frames, we demonstrated once again the validity and efficacy of the Lorentz-invariant approach. To obtain the pion valence light-cone GPDs from the qGPDs, we use NNLO hybrid-scheme matching along with leading renormalon resummation and renormalization group resummation. The use of leading renormalon resummation significantly reduces the perturbative uncertainty in the hybrid renormalization scheme. Furthermore, the use of renormalization group resummation allows us to determine the range of Bjorken-$x$ for which the LaMET approach is reliable for a specific momentum value $P_{z}$. Our results at zero momentum transfer, specifically the valence light-cone PDF results, are in good agreement with the global analyses Barry_2021 within the range $0.2\lesssim x\lesssim 0.8$. The $P_{z}$ dependence of the final light-cone GPD results is significantly reduced compared to the qGPDs, which indicates the effectiveness of the perturbative matching framework. Our lattice calculations provide a detailed three-dimensional imaging of the pion structure, which can be better visualized in terms of impact-parameter-dependent distribution. From these findings, we observe that the effective transverse size of the pion reduces as $x$ increases, a pattern also observed in the proton Cichy:2023dgk . Additionally, the effective size of the pion is consistently smaller than that of the proton for the considered $x$ region. For future studies, several improvements would enhance our understanding: First, increasing statistics and extending $t_{s}$ for larger momenta would enable reliable 3-state fit results. This would allow for a more thorough investigation of excited state contamination. Second, while we acknowledge uncertainties in our preditions for $x<0.2$ and $x>0.8$ regions, incorporating threshold resummation in perturbative matching procedure could yield more precise results in the large-$x$ region. Finally, studying much larger momentum values would provide better insights into higher-twist contributions.
@@ Acknowledgements.
We thank Shohini Bhattacharya, Yushan Su, and Rui Zhang for their valuable communications. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract Nos. DE-SC0012704, DE-AC02-06CH11357, and within the frameworks of Scientific Discovery through Advanced Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond, and under the umbrella of the Quark-Gluon Tomography (QGT) Topical Collaboration with Award DE-SC0023646. HTD is supported by the NSFC under Grant Nos. 12325508, 12293060, 12293064, and the National Key Research and Development Program of China under Grant No. 2022YFA1604900. QS is partially supported by Laboratory Directed Research and Development (LDRD No. 23-051) of BNL and RIKEN-BNL Research Center. SS is supported by the National Science Foundation under CAREER Award PHY-1847893. YZ is partially supported by the 2023 Physical Sciences and Engineering (PSE) Early Investigator Named Award program at Argonne National Laboratory. This research used awards of computer time provided by the U.S. Department of Energy’s INCITE and ALCC programs at the Argonne and the Oak Ridge Leadership Computing Facilities. The Argonne Leadership Computing Facility at Argonne National Laboratory is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC02-06CH11357. The Oak Ridge Leadership Computing Facility at Oak Ridge National Laboratory is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC05-00OR22725. This research also used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. Computations for this work were carried out in part on facilities of the USQCD Collaboration, funded by the Office of Science of the U.S. Department of Energy.
@@ Appendix A Fit results of the two-point and three-point correlation functions
Figure 13: Two examples of the ratio $R^{t}(z=0)/\bar{P}^{t}$ are shown. Left: $P_{z}$ = 0.97 GeV and $-t$ = 0.95 GeV2, right: $P_{z}$ = 1.94 GeV and $-t$ = 1.69 GeV2. The data points represent the lattice results, and the bands with corresponding colors are the fit results of the lattice data. The grey bands show the results of the bare matrix elements.
In this appendix, we present the detailed analyses of the two-point and three-point correlation functions. As mentioned in the main text, by employing the spectral decomposition formula along with 2-state or 3-state fits of the pion two-point correlators Gao:2020ito ; Gao:2021dbh ; Gao:2021xsm ; Gao:2022iex ; Gao:2022vyh , we can extract energy values of both the ground and excited states of the pion, as well as the corresponding amplitudes. We show the results of the ground and the first excited state energy for different momenta in figure 12. We can see that the fit results for $E_{0}$ agree well with the dispersion relation shown by the solid lines, and in most cases, the results for $E_{1}$ also agree with them within errors.
To obtain the matrix elements, we consider the ratio $R^{\mu}$ given by eq. (12) and perform 2-state or 3-state fits. Here, we use the energy levels and amplitudes obtained from the corresponding fits of the two-point functions. In figure 13, we select two specific cases from the data of $z=0$ and $\mu=t$ to show the results of the ratio as $R^{t}/\bar{P}^{t}$. The left one belongs to the case of $P_{z}$ = 0.97 GeV and $-t$ = 0.95 GeV2, while the right one corresponds to the case of $P_{z}$ = 1.94 GeV and $-t$ = 1.69 GeV2, that is, the largest momentum transfer in this work. The lattice results are shown as the data points with different colors representing the different time separations. The bands with corresponding colors denote the two-state fit results of the lattice data, providing a good description of the lattice results. And the grey bands are the extrapolated results of the ratio, that is, the bare matrix elements.
Figure 14: Two examples of the ratio $R^{t}(z=0)/\bar{P}^{t}$ used to study the effectiveness of a 2-state fit with $t_{s}/a\in[9,12,15]$ for large momentum. Left: $P_{z}$ = 1.45 GeV at $-t$ = 0 GeV2, right: $P_{z}$ = 1.94 GeV at $-t$ = 1.10 GeV2. The data points and the bands have the same meaning as in figure 13.
To address excited state contamination for the largest two momenta, we conducted additional analyses of $R^{t}(z=0)/\bar{P}^{t}$ using multiple approaches. Figure 14 presents two representative cases demonstrating the reliability of a 2-state fit with the largest $t_{s}$ being 0.6 fm. For $P_{z}$ = 1.45 GeV at $-t$ = 0 GeV2 (left panel), we compare the results from 2-state (line-filled bands) and 3-state fits (solid-filled bands) using $t_{s}\in[0.36,0.48,0.6]$ fm ($t_{s}/a\in[9,12,15]$). The agreement between these two fits within uncertainties supports the adequacy of 2-state fit at large momentum. For $P_{z}$ = 1.94 GeV at $-t$ = 1.10 GeV2 (right panel), we extended our analysis to larger source-sink separations. We compare 2-state fits using two ranges of $t_{s}\in[0.36,0.48,0.6]$ fm (solid-filled bands, labeled as $t_{s}/a=9-15$) and $t_{s}\in[0.48,0.6,0.72]$ fm (line-filled bands, labeled as $t_{s}/a=12-18$). The agreement between fits across different $t_{s}$ ranges supports the reliability of the ground state extraction. Furthermore, the right sub-figure of the right panel shows the bare matrix element results at $z/a=0$. When we attempted 3-state fit with the extended $t_{s}$ range, the combination of high momentum and large momentum transfer led to prohibitively large uncertainties, making these fits statistically unreliable. The additional analyses above demonstrate that a two-state fit with the largest $t_{s}$ being 0.6 fm can provide stable and reliable results, even at large momenta and momentum transfers.
We collect all the bare matrix element results of the largest momentum in figure 15. It is evident that the results display a decreasing trend with increasing $z$ and also as the momentum transfer increases, as expected.
@@
Appendix B Detailed analysis of scale variation parameter selection
In an OPE, it is natural to choose $\kappa=1$ so that the apparent logarithms vanish in the Wilson coefficients. In practice the choice of $\kappa$ could depend on the observable, but it still should not be too different from $1$. After all, regardless of the exact choice, the key is to estimate the associated theory uncertainty by varying $\kappa$ in a given range, such as $\kappa\in[0.5,2]$ that is typically used in collider phenomenology. In this work, the range of $\kappa$ are mainly decided by the behavior of $C_{0}$ and $\bar{m}_{0}$. Since the studies of $C_{0}$ and $\bar{m}_{0}$ at NNLO+LRR+RGR level are very limited, it is valuable to make a comparison with recent findings from ref. Zhang:2023bxs . In figure 16, we compare our results for $C_{0}(z,\mu=2\text{ GeV})$ (left panel) and $\bar{m}_{0}$ (right panel) derived from NNLO+LRR+RGR perturbative calculations with those from ref. Zhang:2023bxs . Our results are shown in green, while those from the reference are displayed in orange. The results for $\kappa\in[0.75,1,1.5]$ are represented with distinct line styles.
Figure 16: Comparison of $C_{0}(z,\mu=2\text{ GeV})$ (left panel) and $\bar{m}_{0}$ (right panel) derived from NNLO+LRR+RGR perturbative calculations between our results and ref. Zhang:2023bxs .
We observe that at $\kappa=1$ our results and the results of ref. Zhang:2023bxs more or less agree. Both results exhibit larger scale variation when $\kappa<1$. However, for large $z$ the variation with $\kappa$ is larger in our case. Furthermore, the result of ref. Zhang:2023bxs increases for $\kappa<1$, while our result decreases. This discrepancy is due to the differences in how the scale variation parameter ($\kappa$) is fixed: we use a consistent choice of the scale variation parameter throughout the analysis. In ref. Zhang:2023bxs the authors fix the scale variation parameter in the evolution factor before the asymptotic form to 1, while allowing variation in $\kappa$ in other parts of the expression. (More details about the approach used in the reference can be found in a PhD thesis Zhang:2023tnc .) If we modify our formula to match the procedure in ref. Zhang:2023bxs , the corresponding results (shown in purple in figure 16) align reasonably well with those in ref. Zhang:2023bxs .
Figure 17: The results of running coupling constant derived with various values of $\kappa$ using NNLO(NLO)+LRR+RGR perturbative calculations. The solid lines correspond to the NNLO results, while the dashed lines correspond to the NLO results.
Figure 18: Same as figure 2, but with a broader range of $\kappa$.
Regarding the significant scale dependence at lower values of $\kappa$ in our consistent scale variation coefficient approach, we explore a broad range of $\kappa$ to calculate $\alpha_{s}$ and $\bar{m}_{0}$ as shown in figures 17 and 18. Based on these results, at short distances where perturbation theory is applicable, they behave well as expected for $21<\kappa<$ 2, since $\alpha_{s}$ remains reasonably small and $\bar{m}_{0}$ remains constant. However, for lower $\kappa$ values, $\alpha_{s}$ becomes larger and $\bar{m}_{0}$ diverges under the NNLO+LRR+RGR framework. Consequently, we select $\kappa\in[1,1.414,2]$ as our range for scale variation, ensuring the validity of perturbative calculations while maintaining reasonable scale variation to estimate uncertainties.
Appendix C Dependence of the results on $\bar{m}_{0}$ at different values of $z$
To investigate the dependence of the results on $\bar{m}_{0}$ at different values of $z$, we consider the case of $-t=1.095$ GeV2 with NNLO+LRR+RGR($\kappa=1$) perturbative matching as an example, selecting additional values of $\bar{m}_{0}$ at $z/a=4$ and 5 to do the renormalization. The renormalized matrix element $M^{\bar{R}}$ and the quasi-GPD, shown in figure 19, are then compared with those calculated using $\bar{m}_{0}$ at $z/a=3$. It can be found that variations in $\bar{m}_{0}$ across different $z$ values have minimal effect on the renormalized results, as the difference in $\bar{m}_{0}$ between $z/a=3$ and 5 ($\sim$ 0.05 GeV) is negligible compared to the value of $\delta m+\bar{m}_{0}$ ($\sim$ 0.93 GeV). Therefore, using $\bar{m}_{0}$ from different $z$ does not significantly impact our results on the quasi-GPD as can be seen from the right panel of figure 19.
Figure 19: Comparison of the renormalized matrix element $M^{\bar{R}}$ (left panel) and quasi-GPD $\widetilde{H}$ (right panel) derived using $\bar{m}_{0}$ at various values of $z$ for the case of $-t=1.095$ GeV2 with NNLO+LRR+RGR($\kappa=1$) perturbative matching.
Appendix D Extrapolation of the renormalized matrix elements
To test the dependence of extrapolated matrix elements and light-cone GPD results on the fit range, we select four different values of $N\in[3,4,5,6]$, representing the number of data points used in the fit, to perform the fit with the decay model. Taking the case of $-t=1.095$ GeV2 as an example, we compare renormalized matrix elements and light-cone GPD results in figure 20 and 21, respectively. These results are all in good agreement, and the dependence on the fit range is very small. To better follow the extrapolation at long distances, we choose $N=3$ for the analyses throughout this work.
Figure 20: Comparison of extrapolated matrix elements obtained from different fit ranges.
Figure 21: Comparison of light-cone GPD results derived from various fit ranges in the extrapolation step.
Appendix E Discussion of the $-t/P_{z}^{2}$ power correction
In figure 22, we investigate potential power corrections by comparing two data sets with similar values of $-t$ but different $P_{z}$: $-t=1.048$ GeV2 with $P_{z}=1.453$ GeV (yielding $-t/P_{z}^{2}\approx 0.45$ and $-t=1.095$ GeV2 with $P_{z}=1.937$ GeV (yielding $-t/P_{z}^{2}\approx 0.29$). Despite the approximately 50% difference in $-t/P_{z}^{2}$, the results obtained both with and without RGR are comparable between these two cases. This consistency is particularly relevant for our analysis at the largest value of $-t$ (1.69 GeV2), which has a similar $-t/P_{z}^{2}\approx 0.45$, suggesting that the $-t/P_{z}^{2}$ power corrections are well under control even at our largest momentum transfer.
Figure 22: Comparison of LC GPDs with and without RGR.
Appendix F Perturbative coefficient
F.1 Matching coefficient